OpenStudy (anonymous):

Solve the following radical equation (sqrt of 4z+41) +3 = z+2

OpenStudy (kropot72):

\[\sqrt{4z+41}+3=z+2\] Subtracting 3 from both sides gives: \[\sqrt{4z+41}=z-1\] The next step is to square both sides. Can you do that?

OpenStudy (anonymous):

okay i got 4z+41=z^2-2z+1

OpenStudy (kropot72):

Good!. Now subtract 4z and 41 from both sides to get a quadratic of the form \[az ^{2}+bz+c=0\]